OpenStudy (anonymous):

4. Which of these numbers is between 1 and 3/5 (Points:1) A)8/5 B) 5/9 C)5/8 D)3/8

OpenStudy (studygurl14):

3/5 = .6 1= 1 8/5 = 1 3/5 Nope 5/9 = .555555555555555555555555555555555555 Is this between 1 and .6? Nope... Kepp going. :)

OpenStudy (anonymous):

so i just gotta turn them all in to decimals

OpenStudy (anonymous):

what was the answer

OpenStudy (anonymous):

@CassandraMarquez the answer is D, 3/8